2. A string of mass 3.25g is measured to have a length of 3.000m. If a mass of 750g is hung off of this string what is the speed of waves traveling on this string? Show your work
Answer #1

velocity in string under tension

v^2 = T/u

v^2 = mg / u

v^2 = 0.75*9.8 / ( 3.25*10^-3/3)

v = 82.369 m/s

=========

do comment in case any doubt, will reply for sure.. Goodluck